OpenStudy (zeesbrat3):

Water is drained out of tank, shaped as an inverted right circular cone that has a radius of 6cm and a height of 12cm, at the rate of 3 cm3/min. At what rate is the depth of the water changing at the instant when the water in the tank is 9 cm deep? Give an exact answer showing all work and include units in your answer.

OpenStudy (zeesbrat3):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What do you have so far?

OpenStudy (zeesbrat3):

r = 6cm h = 12cm r = 3cm^3/min \[v = \pi r^2h\]

jimthompson5910 (jim_thompson5910):

`r = 3cm^3/min` is incorrect I think you meant to say dV/dt = -3 to represent the fact that the volume is decreasing by 3 cubic cm

jimthompson5910 (jim_thompson5910):

also, you have the wrong volume formula

jimthompson5910 (jim_thompson5910):

volume of a cone \[\Large V = \frac{1}{3}\pi*r^2*h\]

OpenStudy (zeesbrat3):

Oh... So then you take the derivative, yes?

jimthompson5910 (jim_thompson5910):

yeah but first you have to somehow get everything in terms of h you have to find a connection between r and h and do a substitution

jimthompson5910 (jim_thompson5910):

a picture might help 

OpenStudy (zeesbrat3):

r = 1/2h?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

yeah that looks good, r = h/2

jimthompson5910 (jim_thompson5910):

plug that in, derive, then isolate dh/dt

OpenStudy (zeesbrat3):

\[\frac{ 2 }{ 3 } \pi (\frac{ h^2 }{ 2 })\]

jimthompson5910 (jim_thompson5910):

When you plugged r = h/2, and simplified, did you get \[\Large V = \frac{1}{12}\pi h^3\] ??

OpenStudy (zeesbrat3):

So you get \[\frac{ dv }{ dt } = \frac{ 1 }{ 4 } \pi h^2 \frac{ dh }{ dt }\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

dv/dt = -3 h = 9 solve for dh/dt

OpenStudy (zeesbrat3):

\[-3 = \frac{ 1 }{ 4 } \pi (9)^2 \frac{ dh }{ dt }\] \[\frac{ -12 }{ 81 \pi } = \frac{ dh }{ dt }\]

jimthompson5910 (jim_thompson5910):

make sure you reduce the fraction as much as possible

OpenStudy (zeesbrat3):

\[\frac{ -4 }{ 27 \pi } = \frac{ dh }{ dt }\]

jimthompson5910 (jim_thompson5910):

yep \[\Large \frac{dh}{dt} = -\frac{4}{27\pi} \approx -0.047157\] the units for dh/dt are cm/min the rate of the depth of the water is changing at roughly -0.047157 cm/min at exactly the depth of h = 9 cm

OpenStudy (zeesbrat3):

Thank you!!

jimthompson5910 (jim_thompson5910):

glad to be of help